OpenStudy (anonymous):

a how many sided regular polygon has interior angles of 160 degrees? pls show method

OpenStudy (anonymous):

180(n-1)/n =160 180n -180=160n -20n=-180 n=9 sided polygon

OpenStudy (anonymous):

ummm... isn't it n-2?? I think i got it. had a blond moment. i think its 18.